### v1

# The Harmonic Exponential Filter for Nonparametric Estimation on Motion Groups

Abstract
Bayesian estimation is a vital tool in robotics as it allows systems to update the belief of the robot state using incomplete information from noisy sensors. To render the state estimation problem tractable, many systems assume that the motion and measurement noise, as well as the state distribution, are all unimodal and Gaussian. However, there are numerous scenarios and systems that do not comply with these assumptions. Existing non-parametric filters that are used to model multimodal distributions have drawbacks that limit their ability to represent a diverse set of distributions. In this paper, we introduce a novel approach to nonparametric Bayesian filtering to cope with multimodal distributions using harmonic exponential distributions. This approach leverages two key insights of harmonic exponential distributions: a) the product of two distributions can be expressed as the element-wise addition of their log-likelihood Fourier coefficients, and b) the convolution of two distributions can be efficiently computed as the tensor product of their Fourier coefficients. These observations enable the development of an efficient and exact solution to the Bayes filter up to the band limit of a Fourier transform. We demonstrate our filter’s superior performance compared with established nonparametric filtering methods across a range of simulated and real-world localization tasks.  [^1] Project page https://montrealrobotics.ca/hef/
I Introduction
Recursive Bayesian filtering is a powerful approach for state estimation in robotics. It allows the fusion of several different sensor streams while producing belief estimations over the state that can be used in downstream tasks such as planning, control, and mapping. To render the state estimation problem computationally tractable, certain assumptions are frequently invoked, including the presumption that the state distribution, as well as the motion and measurement noise, are unimodal and Gaussian distributed. However, there are many scenarios where a unimodal Gaussian belief fails to place a high probability (density) over the true state. Estimates involving uncertain data associations [1], propagating errors through rigid-body motions [2], and measurements that suffer from multipath errors [3] and symmetries [4] all result in distributions that are distorted or multimodal.
Gaussian filters have been extended to handle some non-Gaussian distributions by modeling the error of the state in Lie groups using parametric distributions [5, 6]. However, to propagate errors, these methods rely on approximations or parametric assumptions to maintain a tractable solution.
((a)) Product
((b)) Convolution
Figure 1: The product (a) and convolution (b) operations between two distributions play an important role in recursive Bayesian filtering. The examples here are shown with harmonic exponential distributions on the circle $S^{1}$.
Beyond Gaussian methods, nonparametric filtering offers an alternative to filters reliant on unimodal distributions or Gaussian assumptions. Examples include particle filters [7, 8, 9] and histogram filters [10, 11, 12]. However, these approaches are not free of limitations. Particle filters are susceptible to particle deprivation [1, Chapter 4], where the correct state is no longer represented within the set of particles. Similarly, histogram filters require resource-intensive computations, including convolutions during the motion propagation phase, and are only accurate up to the quantization error when computing the probability bins.
In this paper, we propose a novel nonparametric and multimodal filtering approach that leverages harmonic exponential distributions [13]. We name this method the harmonic exponential filter (HEF) and it relies on three key benefits of harmonic exponential distributions. First, the prediction step takes advantage of the Fourier analysis convolution theorem, yielding an efficient tensor product instead of a convolution during motion propagation. Second, the update step of the Bayesian filter becomes an element-wise addition. Finally, the Fourier coefficient representation of the harmonic exponential distribution represents more naturally smoothly-varying density functions when compared to histogram binning. These steps are illustrated in Fig 1.
In summary, our main contributions are:
  • An exact approach to computing the posterior belief of the Bayes filter on a compact Lie group, for band-limited prior, motion, and measurement likelihood.
  • A novel nonparametric recursive filtering approach for non-commutative motion groups that can capture multimodal distributions.
  • A filtering strategy whose formulation respects the structure of motion groups and their symmetries.
  • An efficient motion step implementation for motion groups that leverages the Fourier convolution theorem.
We show that our method outperforms the baselines in three non-Gaussian filtering tasks using one simulated and two real-world datasets.
II Related Work
Various approaches have extended Bayesian filtering beyond the unimodal Gaussian assumptions of the Kalman filter (KF). For instance, the KF has been adapted to operate with a mixture of parameterized distributions instead of a unimodal Gaussian [14]. Gilitschenski et al. [15] adapted a mixture model to use direction statistics on the manifold of rotation groups. However, a key challenge with such approaches is the growth of the number of mixture components with each iteration. Common approaches to overcome this challenge are based on distance metrics to merge mixture components [16] or resampling strategies [17].
The KF has also been extended to exploit symmetries of the system by characterizing the state of the robot using Lie groups [18]. If the state error is modeled in the Lie algebra of motion groups with a Gaussian belief and then mapped back to the group with the exponential map, a non-Gaussian posterior is obtained on the manifold [5, 2]. However, to propagate the error state over time and attain a tractable solution, these methods rely on second-order approximations.
Unlike the various KF approaches, our method is nonparametric and can fit any distribution in a broader family of distributions while preserving system symmetries. A well-known nonparametric filtering approach is the histogram filter, which approximates the posterior belief by discretizing the state space and assigning a probability value to each bin [10]. These methods can be adapted to the topology and action of the group [19] and have been widely studied to model distributions on rotation groups [11, 12]. Our method resembles histogram filters as we require discrete samples of the manifold to represent the posterior. However, our formulation operates in both rotation and motion groups and is more computationally efficient during the propagation step than histogram filters. What is more, our approach is exact up to the band limit of a discrete Fourier transform.
Another family of nonparametric and multimodal filters are particle filters which represent the posterior belief using a discrete set of particles [7]. These methods can model general distributions, but are known to suffer from particle deprivation, particularly in high-dimensional state spaces [1, Chapter 4]. Research efforts have been devoted to addressing this problem by proposing better sampling schemes [8, 9], and by changing how particles evolve in the prediction step to better match the posterior distribution[20]. In contrast, our method does not use particles to model the posterior belief and does not suffer from particle deprivation.
Kernel density estimates are a prevalent method to represent nonparametric probability density functions. They can be adapted to handle belief propagation through probabilistic graphical models [21], and the product of distributions within a filtering framework [22].
Pfaff et al. [23] presented an approach that is conceptually similar to ours that they called Fourier filtering. They showed their work on the unit circle and the unit sphere [24, 25]. Although their approach resembles ours in that the probability density functions are represented by Fourier coefficients, our approach extends this concept in several meaningful ways. First, since we model density functions in the log-likelihood space, the product of distributions becomes an element-wise addition instead of the more intricate tensor product needed for Fourier filtering. Second, by representing distributions with the Fourier coefficients derived from the log-likelihood function, the corresponding probability density function is guaranteed to be non-negative. Lastly, our approach can represent densities on locally compact subsets of motion groups $SE(N)$, particularly, any matrix Lie group and not just on a circle or a sphere.
III Harmonic Exponential Filtering
Fourier analysis provides a way to parameterize well-behaved functions in $\mathbb{R}^{N}$ (i.e. one that is square integrable $\mathcal{L}^{2}$ on $\mathcal{V}\subseteq\mathbb{R}^{N}$ and has a finite number of discontinuities in any finite interval in $\mathbb{R}^{N}$) using an infinite sum of trigonometric basis functions. This approach has proven to be useful for various engineering applications for several reasons. Moreover, the convolution theorem shows that computing integrals in spatial domains can be replaced by an efficient tensor product in the frequency domain. A generalized form of Fourier analysis can be used on locally compact subsets of motion groups. The topic is presented in depth by Chirikjian and Kyatkin in [26, Chapters 8,10]. This section presents an overview of our approach. First by providing a brief introduction to Fourier analysis on motion groups (Sec. III-A), then showing how that is used in the family of harmonic exponential distributions to represent a varied set of densities (Sec. III-B), and finally how these concepts can be combined to perform Bayesian filtering (Sec. III-C).
III-A Harmonic Analysis on Locally Compact Motion Groups
Motion groups are a specific category of Lie groups [18] that serve as a representation of rigid-body motions within the Euclidean space $\mathbb{R}^{N}$. This classification includes other important subgroups such as rotation groups $SO(N)$ and affine transformation groups $SE(N)$. More precisely, these are Lie groups because they satisfy the group axioms (see [26, Chapter 7]) and are also a differentiable manifold. Groups are equipped with a set $G$, alongside a binary operator $\circ$, which preserves the closure property, that is, $\forall g,h\in G$, their composition $g\circ h\in G$.
Conducting Fourier analysis on functions supported on Lie groups necessitates a collection of carefully chosen basis functions. These basis functions are in turn the matrix elements of irreducible unitary representations (IURs). A more concrete definition of IURs is provided below.
A representation denotes a mapping $R:G\rightarrow GL(\mathcal{V})$ of a given group $G$ to the set of linear invertible transformations of a vector space $\mathcal{V}$ such that it respects the group operator, $R(g)R(h)=R(g\circ h)$, $\forall g,h\in G$. Here, $g\circ h$ denotes composition of group elements whereas $R(g)R(h)$ is matrix multiplication given an appropriate basis for $\mathcal{V}.$
A unitary representation, $U$, is a representation that does not affect the norm of the vectors on which the representation acts, that is, $\forall g\in G,\forall v\in\mathcal{V}:||U(g)v||=||v||$. Given a unitary matrix $F$, an irreducible unitary representation (IUR) is a unitary representation that cannot be block diagonalized further into another representation via consecutive transformations of the form $F^{-1}U(g)F$. An IUR is represented as $T^{\lambda}(g)\in\mathbb{R}^{m\times n}$ where $\lambda\in\mathbb{Z}^{+}$ is used to index the matrix elements of different IURs.
IURs are the foundation for performing harmonic analysis on groups. If we have a function, $f:G\rightarrow\mathbb{R}$, that is square integrable, and an IUR of group $G$, $T^{\lambda}(G)$, then we can perform a generalized Fourier transform to compute the frequency components of $f$ as
$\mathcal{F}[f]_{mn}^{\lambda}=\eta_{mn}^{\lambda}=\int_{G}f(g)T_{mn}^{\lambda}(g)^{\dagger}d(g),$ (1)
where $d(g)$ is the bi-invariant volume element of $G$ and $(\cdot)^{\dagger}$ the Hermitian transpose. Note that matrix elements of the Fourier coefficients $\eta^{\lambda}\in\mathbb{R}^{m\times n}$, and IUR $T^{\lambda}(g)\in\mathbb{R}^{m\times n}$ are indexed by their row and column indices $m$ and $n$. Using orthonormality between different matrix elements of $T(g)$ [13], the generalized inverse Fourier transform is
$\mathcal{F}^{-1}[\eta](g)=f(g)$ $=\sum_{\lambda}\sum_{n,m}\eta_{mn}^{\lambda}T_{nm}^{\lambda}(g).$ (2)
For the work presented in this paper, we use the IUR of $SE(2)$ presented in [26, 27]. For $g(x,y,\theta)\in SE(2)$ with $(x,y)\in\mathbb{R}^{2}$ and $\theta\in[0,2\pi]$ parameterizing an element of $SE(2)$, a variable $\psi\in[0,2\pi)$, bi-invariant volume $d\mu\triangleq dx\,dy\,d\theta\,d\psi$ and complex unit $i$, the Fourier coefficient $\eta^{\lambda}_{mn}$ can be computed as
[…]
Chirikjian and Kyatkin provide a thorough description of the computation of the fast Fourier transform (FFT) over 2D motion groups in [26, Section 11.2]. Henceforth, we will omit indices and denote the generalized Fourier transform for (locally) compact motion groups as $\eta=\mathcal{F}[f]$.
In this work, the FFT is used to compute the Fourier coefficients from a set of discrete samples of the function on the group. In practice, to avoid an infinite sum or integration, the Fourier analysis needs to be performed on a locally compact subset of a group. For $SE(2)$, the orientation portion, $SO(2)$, is already compact. For translation, we only consider a predefined locally compact subset of $\mathbb{R}^{2}$, as it is presented in [26].
III-B Harmonic Exponential Distributions
[…]
$p_{\eta}(g)\triangleq p(g\mid\eta)=\frac{1}{Z_{\eta}}\text{exp}(\eta\cdot T(g)),$ (3)
where $\eta$ are the parameters, $T(g)$ the IUR of group $G$, $Z_{\eta}$ the normalizing constant, and ()($\cdot$)( ⋅ ) is an appropriate inner product. For example, $\forall g\in SO(2)$ their IURs are $1\times 1$ matrices of the form $T_{00}^{\lambda}(g)=e^{i\lambda g}$, hence, both $T(g)$ and $\eta$ are vectors whose inner product ()($\cdot$)( ⋅ ) is the standard dot product. Note that this is equivalent to computing (2) on $SO(2)$. The parameters $\eta$ and $Z_{\eta}$ can be efficiently computed via Harmonic analysis from $\phi(g\mid\mathbf{\eta})\triangleq\exp(\mathbf{\eta}\cdot T(g))$ as
$\begin{split}\ln\phi(g\mid\eta)&=\eta\cdot T(g)=\mathcal{F}^{-1}[\eta](g),\\ Z_{\eta}&=\int_{G}\phi(g\mid\eta)T^{0}_{00}(g)d(g)=\mathcal{F}[\phi]^{0}_{00},\end{split}$ (4)
with $\mathcal{F}[\phi]^{0}_{00}$ being the zeroth Fourier coefficient and $T^{0}_{00}(g)=1$. Representing the density with Fourier coefficients in log-likelihood space, as opposed to directly representing the density, ensures the probability density function is non-negative for any set of parameters $\eta$, and the normalizing factor $Z_{\eta}$ ensures that it integrates to 1. From this point forward, we use the shorthand notation $p_{a}(g)\triangleq p_{\eta_{a}}(g)$. The harmonic exponential distribution provides an elegant solution for the product of two distributions, achieved through the element-wise addition of their Fourier coefficients:
$p_{a}(g)p_{b}(g)$ $=\frac{1}{Z_{\eta}^{a}Z_{\eta}^{b}}\exp((\eta_{a}+\eta_{b})\cdot T(g))$ (5)
$=p(g|\eta_{a}+\eta_{b}).$
Figure 2: Relationship between probability $p(g)$, log-likelihood $\text{ln}(p(g))$, and their respective spectral coefficients $\eta$ and $M$. We model the log-likelihood $\text{ln}(p(g))$ using its spectral coefficients $\eta$. To recover the density $p(g)$, we compute the inverse FFT of $\eta$ and exponentiate the result. The spectral coefficients $M$ are computed by taking the FFT of $p(g)$ and are used to compute convolutions.
Convolving two distributions is inherently more complicated. However, harmonic analysis provides a connection between convolutions and spectral coefficients,
$(p_{a}\ast p_{b})(g)$ $=\int_{G}p_{a}(h)p_{b}(h^{-1}\circ g)d(h)$ (6)
$=\mathcal{F}^{-1}[\mathcal{F}[p_{b}]\mathcal{F}[p_{a}]](g)$
$=\mathcal{F}^{-1}[M_{b}M_{a}](g),$
with $M_{a}=\int_{G}p_{a}(g)T(g)d(g)=\mathcal{F}[p_{a}]$ being the Fourier coefficients of the probability density function $p_{a}$, instead of a log-likelihood function as in (4). Due to the nonconmutative nature of $SE(N)$, encoded through $T(g)$, the order of the Fourier coefficients is reversed in (6). Observe that $M_{a}$ and $M_{b}$ are composed via matrix products: $M^{\lambda}=M_{b}^{\lambda}M_{a}^{\lambda}$, $\forall\lambda\in\mathbb{Z}^{+}$. A complete proof is presented by Chirikjian and Kyatkin [26, Section 10.4]. The relation between these spaces can be seen in Fig. 2.
III-C Bayesian Filtering
Following the notation presented by Thrun et al. [1, Chapter 2], equations (5) and (6) allow us to incorporate the harmonic exponential distribution into the Bayesian filtering framework:
$bel(x_{t})\triangleq p(x_{t}|z_{1:t},u_{1:t})$ $=\frac{p(z_{t}|x_{t})p(x_{t}|z_{1:t-1},u_{1:t})}{p(z_{t}|z_{1:t-1},u_{1:t})}$ (7)
$=\gamma p(z_{t}|x_{t})p(x_{t}|z_{1:t-1},u_{1:t})$
$=\gamma p(z_{t}|x_{t})\overline{bel}(x_{t}),$
where the predicted belief over the state $x_{t}$ after incorporating the motion model but before incorporating the measurement $z_{t}$ is denoted as $\overline{bel}(x_{t})\triangleq p(x_{t}|z_{1:t-1},u_{1:t})$, and the posterior belief after incorporating $z_{t}$ as $bel(x_{t})$.
Computing the predicted belief is commonly referred to as the “prediction step” and requires one to define the motion model. The motion model is the likelihood of the current pose $x_{t}\in SE(N)$ given the previous pose $x_{t-1}\in SE(N)$ and control input $u_{t}$, that is, $p(x_{t}|u_{t},x_{t-1})$. If we assume that the motion is fully defined by the control input, the motion model becomes the likelihood of the relative motion from $x_{t-1}$ to $x_{t}$, or $x^{-1}_{t-1}\circ x_{t}$ matching the relative motion induced by $u_{t}$ or $p_{u_{t}}(x^{-1}_{t-1}\circ x_{t})$. Our prediction step becomes
[…]
$=\mathcal{F}^{-1}[\mathcal{F}[p_{u_{t}}]\mathcal{F}[bel_{t-1}]](x_{t}).$
Algorithm 1 The Harmonic Exponential Bayes Filter
Input: Previous posterior belief $bel_{t-1}$, relative motion likelihood $p_{u_{t}}$, current measurement likelihood $p_{z_{t}|x_{t}}$ and IUR $T$.
1:// Prediction step
2:$\overline{bel}(x_{t})=\mathcal{F}^{-1}[\mathcal{F}[p_{u_{t}}]\mathcal{F}[bel_{t-1}]](x_{t})$triangleright Eq. (8)
3:// Fourier coefficients
4:$\eta_{z_{t}|x_{t}}\leftarrow\mathcal{F}[\ln(p_{z_{t}|x_{t}})]$triangleright Eq. (4)
5:$\eta_{b\overline{el}_{t}}\leftarrow\mathcal{F}[\ln(\overline{bel_{t}})]$triangleright Eq. (4)
6:// Update step
7:$bel(x_{t})=\gamma\exp((\eta_{z_{t}|x_{t}}+\eta_{b\overline{el}_{t}})\cdot T(x_{t}))$triangleright Eq. 9
8:return $bel_{t}$
The update step of a Bayes’ filter is the product of two probability distributions, $p(z_{t}|x_{t})\overline{bel}(x_{t})$. If we perform Fourier analysis in the domain $x_{t}$ of the measurement log-likelihood function we can get a set of Fourier coefficients $\eta_{z_{t}|x_{t}}\triangleq\mathcal{F}[\text{ln}(p(z_{t}|x_{t})]$, and the measurement likelihood function becomes $p(z_{t}|x_{t})=\frac{1}{Z_{\eta}}\exp(\eta_{z_{t}|x_{t}}\cdot T(x_{t}))$. With that and using (5), the update step simplifies to
$bel(x_{t})=\gamma\exp((\eta_{z_{t}|x_{t}}+\eta_{b\overline{el}_{t}})\cdot T(x_{t})),$ (9)
where the normalizing constant $\gamma$ is computed from the zeroth Fourier coefficient as shown in (4).
From this point onward, we use the notation $bel_{t}:SE(N)\to\mathbb{R}^{\geq 0}$ to denote the probability density function at time $t$ associated with the posterior distribution. The predicted belief $\overline{bel}_{t}$, motion model $p_{u_{t}}$, and measurement likelihood $p_{z_{t}|x_{t}}$ are denoted similarly. Note that the previous definitions are functions and not point estimates. Therefore, we can compute the Fourier coefficients of our density functions as $\eta_{z_{t}|x_{t}}\triangleq\mathcal{F}[\ln(p_{z_{t}|x_{t}})]$ and $\eta_{b\overline{el}_{t}}\triangleq\mathcal{F}[\ln(\overline{bel_{t}})]$. A detailed solution to the Bayes’ filter using the harmonic exponential distributions is presented in Algorithm 1.
III-D Analysis of the Harmonic Exponential Filter
We would now like to point out several advantages inherent to the formulation of the HEF. One is the efficiency compared to a histogram-based approach. The HEF requires a transition between two different Fourier spaces to calculate the convolution result, which inherently adds complexity. Nevertheless, it can still be more efficient than a standard convolution implementation in 3D. If we consider a discretized space of $n\times n\times n$ in 3D, a convolution has complexity $O(n^{6})$, while a 3D FFT is of $O(n^{3}\log^{3}{n})$, and the tensor multiplication needed for (6) would be $O(n^{4})$.
Convolution Runtime (s)
Grid size $(10,10,8)$ Grid size $(20,20,8)$ Grid size $(40,40,8)$
Hist (PP) 0.2284 15.7969 997.2984
Hist (Scipy) 0.0375 2.2434 146.4707
HEF (Ours) 0.0014 0.0083 0.0684
TABLE I: A runtime comparison between a histogram-based approach and the HEF for computing the convolution of two functions. Hist (PP) is a direct convolution written in pure python. Hist (Scipy) uses the scipy.signal.convolve() function, and HEF (ours) is the approach we presented in (6) implemented in python using the Numpy FFT library.
Table I shows a runtime comparison between a histogram-based approach and the HEF over varying sample grid sizes. As shown, the histogram approach quickly becomes intractable. A workaround we use for the benchmark algorithm in Sec. IV is to convolve with a smaller kernel, eliminating areas with near zero probability, but that is only an option for certain densities, such as low covariance Gaussians.
((a))
((b))
Figure 3: Analysis of the harmonic exponential distribution and histogram distributions in modeling a mixture of two von Mises distributions on $S^{1}$. (a) depicts how each method reproduces the underlying mixture. (b) shows how $D_{KL}$ evolves for each method as the number of parameters used increases, and shows the effect of the concentration parameter $\kappa$ on the fidelity of each method, as measured by $D_{KL}$.
Another advantage of representing the density function with Fourier coefficients instead of the probability bins used in histogram filters is the fidelity with which they can represent signals. When computing the Fourier coefficients for the harmonic exponential distribution, the FFT operations behave as a low-pass filter on the signals, leading to information loss in the distribution. However, binning a probability distribution also causes a loss of information. To compare the two, we computed the Kullback–Leibler (KL) divergence $D_{KL}(P||Q)$ between the two approaches and a mixture of two von Mises distributions on $S^{1}$. This analysis is shown in Fig. 3. The KL divergence is a measure of the distance between two distributions, computed as
$D_{KL}(P||Q)=\int_{\mathcal{X}}p(x)\log\left(\frac{p(x)}{q(x)}\right)dx.$
The von Mises distribution is a probability distribution on the circle $S^{1}$ whose density function is given by
$p(g|\mu,\kappa)=\frac{\exp\left(\kappa\cos(x-\mu)\right)}{2\pi I_{0}(\kappa)},$
((a))
((b))
Figure 4: Ability of the HEF and other classic filtering approaches to represent the banana distribution. (a) shows the prior and motion model, which are modelled with a rectangular and Gaussian distribution, respectively. Note that the EKF also uses a Gaussian prior. (b) presents the posterior distribution of the four different approaches after five prediction steps. Only the HEF and particle filter are able to capture the banana distribution.
where $I_{0}(\cdot)$ is the Bessel function and $\kappa$ plays a role similar to the inverse of the covariance in a Gaussian distribution. We can also define a mixture of two von Mises distributions, that is, $p(g)=\frac{1}{2}p(g|\mu_{1},\kappa_{1})+\frac{1}{2}p(g|\mu_{2},\kappa_{2})$. Our analysis in Fig. 3 shows that harmonic exponential distributions have a smaller $D_{KL}$ across a range of situations. When both approaches use few parameters, the harmonic exponential distribution does a better job of representing the underlying density. Increasing the number of parameters improves the fidelity of both approaches; however, the harmonic exponential distribution maintains its advantage in terms of $D_{KL}$. We can also use the $\kappa$ parameter to show the “low-pass filter” effect of taking the FFT. As $\kappa$ increases, the distribution includes more high-frequency components, and $D_{KL}$ increases, but as Fig. 3 shows, the histogram approach performs even worse.
Another benefit of the HEF is its expressivity in capturing diverse distributions, such as the banana shape distribution first presented by Thurn et al. [1, Chapter 5]. This distribution appears when the uncertainty in the heading of an agent increases due to the accumulation of error during multiple prediction steps. Unimodal approaches such as the EKF fail to capture this distribution as it can only model the uncertainty using error ellipses, which is inconsistent with the density of the banana distribution, as shown in Fig. 4.
We compare the expressivity of our method against three other well-established filtering approaches by taking five prediction steps in a horizontal straight path to the right. Interestingly, only the HEF and particle filter are able to capture the banana distribution. Note that all filters start from a rectangular prior, as shown in Fig. 4(a), except for the EKF due to its Gaussian assumptions. Although histogram filters are highly expressive, they fail to capture the banana distribution, as 3D convolutions fail to model the effect of rotations in $SE(2)$.
IV Evaluation
To evaluate our method, we use one simulated and two real-world datasets where the robot state is represented in $SE(2)$. Our assessment focuses on two main metrics. The first metric is the absolute trajectory error (ATE), calculated using two different estimators of the robot position: the mode and mean of the posterior belief. The ATE, measured in meters, is computed along the entire trajectory. However, as this metric does not fully capture the non-Gaussian nature of multimodal estimators, we also calculate the negative log-likelihood (NLL) of the ground truth position under the current belief. We use the equation $\text{NLL}=-\frac{1}{T}\sum_{t}\log bel(x_{t}^{GT})$ where $T$ is the total number of time steps and $bel(x_{t}^{GT})$ is the posterior belief of the true state position, $x_{t}^{GT}$, computed at time $t$. The result of the HEF is the posterior belief over the state space which can be used in a state-belief planner [28]. NLL helps determine how useful our approach would be when used with these systems.
To compute the FFTs, we define a grid of size $50\times 50\times 32$ in $SE(2)$. This grid consists of $50\times 50$ equally spaced Cartesian samples over the interval $(x,y)\in[-0.5,0.5]^{2}$ and 32 equally spaced angular samples $\theta\in[0,2\pi)$. We transform the map of all experiments to the size of this grid. A standard procedure to compute the mean for a harmonic exponential distribution is to take the gradient of the log partition function $\nabla_{\eta}\log Z_{\eta}=\mathbb{E}\left[T(x)\right]\triangleq\mu\quad$ [29, Section 3.4.3]. However, an analytical expression for $\nabla_{\eta}\log Z_{\eta}$ is difficult to compute since it involves solving an integral. We approximate the mean relying on the same samples used to compute the FFTs, that is, $\mu\approx\sum_{n}bel(x_{n})T(x_{n})$. The mode of the HEF is also approximated using the FFT samples: $x^{*}=\underset{x\in{x_{1},x_{2},\dots,x_{N}}}{\operatorname*{argmax}}\mathcal{F}^{-1}[\eta_{bel}](x)$.
In these assessments, our proposed HEF presented in Sec. III is compared against the following filtering approaches: an EKF implementation that assumes unimodal Gaussian state and error, a histogram filter (HistF), and a particle filter (PF). For the sake of comparison, HEF and HistF use $80.000$ parameters, that is, a $50\times 50\times 32$ grid and PF uses $80.000$ particles. To determine the NLL of the PF we approximate the density function using a histogram whose resolution is the same one used for the HEF and HistF. The mode of the PF is estimated by taking the heaviest weighted particle before resampling. For all our experiments we use as a motion model the one of a differential drive robot corrupted with Gaussian noise. To estimate the motion and measurement noise of each filter, we sweep over different noise values and took those that minimize the NLL.
IV-A Simulation
The first evaluation is performed in a simulation environment with range-only measurements. This environment includes a series of landmarks arranged in a line. The robot travels in a counterclockwise circle around the landmarks. The landmarks alternate in providing a range measurement at each time step, where we assume known correspondence. We chose to simulate range measurements due to their highly non-Gaussian measurement likelihood, which produces a ring around each landmark, as can be seen in Fig. 5. We present the results of this experiment in Table II, where metrics are computed over 10 different random seeds. Overall, our method outperforms the baselines, obtaining the lowest ATE and NLL in this multimodal localization experiment.
((a))
((b))
Figure 5: Sample results from the simulated environment. (a) shows the posterior belief for the baselines and the HEF. (b) depicts the steps of the HEF: the predicted belief (top) is $\overline{bel}_{t}$ from (8), the range-only measurement likelihood (middle) is $p_{z_{t}|x_{t}}$, and the posterior belief (bottom) is $bel_{t}$ from (9).
ATE w/ mode (m) ATE w/ mean (m) NLL (-)
EKF 0.093 $\pm$ 0.04 0.093 $\pm$ 0.04 228.23 $\pm$ 27.78
HistF 0.114 $\pm$ 0.03 0.102 $\pm$ 0.02 13.407 $\bm{\pm}$ 0.49
PF 0.138 $\pm$ 0.02 0.101 $\pm$ 0.02 4.040 $\pm$ 0.92
HEF (ours) 0.058 $\pm$ 0.02 0.056 $\bm{\pm}$ 0.02 0.906 $\pm$ 0.90
TABLE II: Quantitative results for the range-only simulator. Reported metrics include the absolute trajectory error (ATE) computed using the mode and the mean of the posterior belief, and the negative log-likelihood (NLL). Lower values are better. We report the mean and standard deviation of these metrics over 10 different random seeds.
IV-B UWB Range Dataset
((a)) UWB.
((b)) Doors.
Figure 6: Setup for the UWB and doors datasets. (a) shows the layout used in the two UWB runs with a Clearpath Husky and five ultra-wideband (UWB) beacons. (b) presents a top-down view of the laboratory used in the doors dataset.
For our first real-world evaluation, we used the Clearpath Husky unmanned ground vehicle (UGV) shown in Fig. 6(a). The UGV is equipped with an ultra-wideband (UWB) tag that communicates with stationary UWB beacons in the surrounding room. Using a ranging protocol between the UGV-mounted UWB tag and the stationary UWB beacons, range measurements between each are computed. The experiment consists of two runs, each with five beacons that were mapped prior to the experiment. For each run, we removed beacons one by one ignoring their range measurements, producing five trials per run. We assume known correspondence.
((a)) Five beacons.
((b)) Three beacons.
((c)) One beacon.
Figure 7: Top-down view showing the configuration of the beacons used in the UWB experiments. By reducing the number of beacons moving from (a) to (c), the measurement likelihood transitions from Gaussian to highly non-Gaussian.
The motion of the agent was determined by integrating the inertial measurement unit (IMU) and the wheel encoders of the Husky platform. The motion noise was estimated to have a standard deviation of 3.08 cm and 0.57 degrees per step. Finally, the ground truth was provided by a NaturalPoint OptiTrack motion capture system. We repeat each run with a different number of beacons to demonstrate the robustness of our proposed HEF to handle unimodal and multimodal distributions. The layout of the beacons used in the experiments is shown in Fig. 7.
The results of this evaluation are presented in Table III. In all runs and with different beacon configurations, the HEF is comparable to or better than the baseline filters even in the presence of highly multimodal measurement likelihoods. Although the EKF obtained the best ATE among all filters in the first run with one beacon, its NLL is the highest among all filters, indicating that throughout the experiment the filter is highly uncertain about the ground truth position. As all the results of the second run closely resembled those of the first run, they were not included in the table.
5 Beacons 1 Beacon
ATE w/ mode (m) ATE w/ mean (m) NLL (-) ATE w/ mode (m) ATE w/ mean (m) NLL (-)
EKF 60161 60161 861e8 0.903 0.903 11.030
HistF 0.221 0.210 -2.653 1.157 1.202 -0.364
PF 0.304 0.210 3.311 3.731 2.639 9.253
HEF (ours) 0.199 0.194 -2.984 1.616 1.076 -0.533
TABLE III: Quantitative results for the UWB dataset, run II. Reported metrics include the absolute trajectory error (ATE) computed using the mode and the mean of the posterior belief, and the negative log-likelihood (NLL). Lower values are better. Note that the EKF is highly uncertain about the ground truth position of the agent during the experiments.  [^3] Complete results at https://montrealrobotics.ca/hef/.
IV-C Doors Dataset
In our final real-world experiment, we draw inspiration from the semantic localization problem discussed in Thurn et al. [1, Chapter 1]. The experiment involves localizing an agent navigating anticlockwise in a corridor using doors detected from RGB images as landmarks. All doors share similar semantic features, such as color or handle. Therefore, we do not assume known correspondence for which of the 21 doors generated a given measurement in this problem.
We premapped the environment using gmapping [30] and assumed a bearing-only measurement model that enforces a multimodal distribution. Bearing measurements are computed using the centroid of the doors. We only compute bearing measurements for samples in the free space of our occupancy grid map. Similar to the experiment in Sec. IV-B, we integrate the IMU and wheel encoders to estimate the motion of a Jackal platform. The motion noise was estimated to have a standard deviation of 3.51 cm and 2.50 degrees per step. The ground truth position was estimated using the laser-based SLAM implementation of gmapping. The estimated entropy of the posterior distribution over the ground truth’s pose using gmapping was $2.232$ nats. The layout of the experiment, as well as an image of a door within the dataset is shown in Fig. 6(b).
ATE w/ mode (m) ATE w/ mean (m) NLL (-)
EKF 2.224 $\pm$ 1.91 2.224 $\pm$ 1.91 521.479
HistF 9.739 $\pm$ 7.78 8.806 $\pm$ 7.12 16.719
PF 5.660 $\pm$ 4.78 3.541 $\pm$ 2.82 4.605
HEF (ours) 1.525 $\pm$ 1.34 1.559 $\bm{\pm}$ 1.36 0.986
TABLE IV: Quantitative results for the doors dataset. Reported metrics include the absolute trajectory error (ATE) computed using the mode and the mean of the posterior belief, and the negative log-likelihood (NLL). We also report the standard deviation of the translation error.
As there is no known correspondence, we employ a greedy data association strategy and compute the measurement likelihood using only the door that is more likely to have generated a given measurement. We present the results of this experiment in Table IV. Our HEF surpasses by a large margin all baselines in this challenging task, showcasing its ability to handle multimodal distributions and the importance of symmetries for robot localization. A snapshot of the filters is shown in Fig. 8.
During these experiments, our proposed approach took an average of 1.14s to run for each timestep. Our approach was implemented in python and ran on an i7-12700H CPU with 16GB of RAM. This was fast enough to be real-time for the door dataset, because the robot would never observe two doors within that time frame.
((a))
((b))
Figure 8: Sample results from the doors dataset. (a) shows the posterior belief for the baselines and the HEF. (b) features the steps of the HEF: the predicted belief (top) is $\overline{bel}_{t}$ from (8), the bearing-only measurement likelihood (middle) is $p_{z_{t}|x_{t}}$, and the posterior belief (bottom) is $bel_{t}$ from (9).
V Conclusion
In this paper, we have presented a novel approach to nonparametric Bayesian filtering using the harmonic exponential distribution. We showed how our method was capable of capturing multimodal distributions while preserving the structure of non-commutative motion groups by relying on harmonic analysis [26]. Harmonic analysis on groups and homogeneous spaces has attracted new interest due to the growing field of geometric deep learning. As work in this field shows, FFT operations and related convolution computations are fully differentiable [31]. This leads to the possibility of using the harmonic exponential filter as a fully differentiable filter for end-to-end training of measurement and motion models. Another promising avenue for exploration is to use harmonic exponential distributions in a SLAM framework or a state-belief planner [28]. As Fourie et al. [32] showed, all that is needed to perform belief propagation in a Bayes tree is the product operation between two distributions.